OpenStudy (narissa):

what point slope form that passes (2,-1) and has a slope of 3.

OpenStudy (narissa):

@imqwerty

imqwerty (imqwerty):

again use this-> :) \(\Huge (y-y_1)=m(x-x_1)\) here the given point is \(x_1,y_1\) and slope is \(m\)

OpenStudy (narissa):

(y-1)=3(x+2)??

imqwerty (imqwerty):

:) you did the same mistake like u did last time we substitute the alues we get this- \(\large [y-(-1)]=3[x-(2)]\) now solve

OpenStudy (narissa):

umm lol im confused is it (y+1)=3(x-2)?

TheSmartOne (thesmartone):

yea

imqwerty (imqwerty):

yes it is this ^

imqwerty (imqwerty):

where are you confused tho? :)

OpenStudy (narissa):

oh okay thank you lol i need to write this down

OpenStudy (narissa):

oh i was confused about where the subtraction and addition signs go. but i get it now